# Data-Dependent Generalization Bounds for Parameterized Quantum Models Under Noise

### v2

@@ 1 Introduction
Quantum machine learning (QML) holds the promise of revolutionizing data analysis by leveraging quantum mechanical principles to enhance computational efficiency and model expressivity [1, 2, 3]. Although still in its early stages [4, 5, 6], QML has shown potential advantages over classical machine learning in certain tasks, such as classification [7, 8], regression [9], and clustering [10]. A crucial aspect of any machine learning model, including QML models, is its generalization ability. Generalization refers to a model’s ability to perform well on unseen data [11, 12, 13]. Generalization bounds provide theoretical guarantees on the performance of a model in new data and play a vital role in ensuring the reliability and robustness of machine learning models [14, 15, 16].
@@
One of the prominent questions in QML is understanding how noise affects the generalizability of the models [17, 18, 19]. Generalization bounds are typically derived by bounding a complexity measure of the hypothesis class used for learning [19]. These complexity measures, such as Rademacher complexity, VC dimension, and metric entropy, quantify the expressiveness or capacity of the model class. Generalization bounds are particularly important due to the inherent challenges posed by quantum systems, especially in the Noisy Intermediate-Scale Quantum (NISQ) era [20, 21, 6, 19, 22, 18]. NISQ devices are characterized by limited qubit coherence times, gate fidelities, and susceptibility to noise, which can significantly impact the performance and reliability of QML models. Several studies have investigated generalization bounds for QML [21, 23, 17, 6, 15, 22, 18, 19, 24, 25, 26], exploring various approaches such as uniform bounds, margin-based bounds, and stability analysis [27]. These studies have provided valuable insights into the factors influencing the generalization capabilities of QML models, including the architecture of quantum neural networks (QNNs), data encoding strategies, and the choice of optimization algorithms. Furthermore, research has explored the impact of noise on QML [4, 28, 29, 30], highlighting the challenges of maintaining model performance in the presence of noise. These noises, arising from various sources such as decoherence [31], gate errors, and other quantum phenomena, pose a formidable challenge in developing QML models that can reliably learn and generalize from the data. Although strategies for mitigating noise [32, 33, 34, 33], have been proposed to improve the performance of QML models, most of the existing research [35, 36, 19, 21, 37, 17, 38, 18] on QML generalization bounds focuses on idealized and noise-free quantum computers. Researchers have also explored the QML model’s generalization capability in the context of quantum kernels [39, 37, 40]. However, these bounds are often tailored to the specific algorithmic technique or the problem at hand. Hence, a comprehensive understanding of how noise affects the generalization performance of QML models is still lacking. This gap is particularly critical in the NISQ era, where noise is unavoidable and can severely limit the practical applicability of QML algorithms. Some studies [29, 41, 4, 36, 40, 30] suggest that quantum noisy channels can significantly hinder the capabilities of QML, potentially affecting their practical applications. We note that the generalization capabilities of quantum neural network models may suffer from the barren plateau problem [42, 43, 44, 45, 29, 46, 47] and the quantum kernel methods from exponential concentration [41]. Please note that the training-induced noise differs from the inherent noise in the quantum devices and is not considered in this work.

### v3

@@ 1 Introduction
Quantum machine learning (QML) holds the promise of revolutionizing data analysis by leveraging quantum mechanical principles to enhance computational efficiency and model expressivity [1, 2, 3]. Although still in its early stages [4, 5, 6], QML has shown potential advantages over classical machine learning in certain tasks, such as classification [7, 8, 9], regression [10], and clustering [11]. A crucial aspect of any machine learning model, including QML models, is its generalization ability. Generalization refers to a model’s ability to perform well on unseen data [12, 13, 14]. Generalization bounds provide theoretical guarantees on the performance of a model in new data and play a vital role in ensuring the reliability and robustness of machine learning models [15, 16, 17].
@@
One of the prominent questions in quantum machine learning is understanding how noise affects the generalizability of models [18, 19, 20]. Generalization bounds are typically derived by bounding a complexity measure of the hypothesis class used for learning [20]; such measures—like Rademacher complexity, VC dimension, and metric entropy—quantify the expressiveness or capacity of the model class. These bounds are especially crucial in the NISQ era [21, 22, 6, 20, 23, 19], where devices are characterized by limited qubit coherence times, suboptimal gate fidelities, and a high susceptibility to noise, all of which can significantly impact the performance and reliability of QML models. Several studies have investigated QML generalization bounds—exploring approaches such as uniform bounds, margin-based bounds, and stability analysis [22, 24, 18, 6, 16, 23, 19, 20, 25, 26, 27, 28, 29]. Recent works by Bu et al.[29] have further refined this framework by rigorously linking the complexity of QML models to their quantum resource content through tailored Rademacher and Gaussian analyses. In particular, they introduce a $(p,q)$-norm to quantify the non-stabilizer (or “magic”) resource, demonstrating that the incorporation of non-Clifford gates—characterized via free robustness measures—directly influences the generalization capacity of quantum circuits[28]. These contributions offer valuable insights into how circuit architecture, data encoding strategies, and optimization choices affect QML generalization. Concurrently, research has also examined the broader impact of noise on QML [4, 30, 31, 32], emphasizing that noise from sources such as decoherence [33], gate errors, and other quantum phenomena poses formidable challenges to developing models that can reliably learn and generalize. Although several strategies for mitigating noise have been proposed [34, 35, 36, 35], much of the existing research on QML generalization bounds focuses on idealized, noise-free quantum computers. Moreover, while the generalization capabilities of quantum neural networks and quantum kernel methods have been explored [37, 38, 39], these bounds are often tailored to specific algorithmic techniques or problems. Some studies [31, 40, 4, 41, 39, 32] further suggest that quantum noisy channels can significantly hinder QML capabilities. We note that the generalization abilities of quantum neural network models may also be compromised by the barren plateau phenomenon [42, 43, 44, 45, 31, 46, 47] and that quantum kernel methods can suffer from exponential concentration [40]. It should be noted that training-induced noise differs from the inherent noise in quantum devices and is not considered in this work. Consequently, a comprehensive understanding of how noise affects the generalization performance of QML models remains lacking—a gap that is particularly critical in the NISQ era, where noise is unavoidable and can severely limit the practical applicability of QML algorithms.